OpenStudy (anonymous):

multiply the polynomials using the distributive property and combine like terms: (x+7)(-3x+2)

OpenStudy (anonymous):

First, multiply x by everything in the second parenthesis, aka "distribute" it. Then do the same with the 7. I trust you know how to combine the like terms after you've done that.

OpenStudy (anonymous):

x*-3x=-3x^2 x*2=2x 7*-3x=-21x 7*2=14 -3x^2+2x-21x+14\[-3x^2-19x+14\]